# SimCity: Multi-Agent Urban Development Simulation with Rich Interactions

### v1

@@ E.3 Full Examples
reasoning’: The city reports significant demand and lack in several key sectors, particularly in Food_Beverages, Basic_Metals, Other_Nonmetal_Mineral_Products, Chemical_Products, and Accommodation_Food_Services, indicating an opportunity for focused production to meet these demands. The layout reveals available building spaces that are not immediately adjacent to residential areas but close enough for workforce accessibility. Given the high lack of Food_Beverages and Basic Metals, building a Food Processing Plant and a Metal Fabrication Plant would be appropriate. The positions chosen in the layout are close enough to existing infrastructure to leverage utilities but are not directly next to residential zones to avoid any negative impacts.’,
@@
{’x’: 12, y’: 2, type’: Food Processing Plant’},
{’x’: 14, y’: 2, type’: Metal Fabrication Plant’}
@@
Here is the monthly report for Educational Institution (No. 200025):
Overall, during this month, the building gross profit is 1040.80, with rest cash 5784.46. It has 1329.2468 durable investments for promoting production.
The building has 0 employees, 6 vacant positions, and run with efficiency 0.0000%.
@@
This month, the building income is 1040.800, which consists of the following parts:
- Local trading income: 1040.800(100.0000%)
Theres no outcome this month
The efficiency of this building is 0.0000%, which consists of the following parts:
- Principal has 28.5714% importance, but no one is working here. If an employee with average skill(Management) works here, about 1804.5412 goods will be produced per month.
- Teacher has 22.8571% importance, but no one is working here. If an employee with average skill(Teaching) works here, about 1543.5823 goods will be produced per month.
[... and other employees]
The building has 6 positions, and here are the details:
- position_id: 2000067, None(Principal), salary: 100
- position_id: 2000068, None(Teacher), salary: 100
[... and other positions]
The building has no employees, so you may consider to increase salary of the position to attract more employees.
The building run with profit: 1040.8000. According to current dividend rate, 50.0000% of current balance will be distributed to shareholders, which is 2892.2278, and the rest 2892.2278 will stay in the building.
The social average return on investment is 10.6795%, and the interest rate of the bank is 37.9503%.
@@
- Education: 24230.216
- Public_Administration_Defence_SocialSecurity: 5.984
- Health_Social_Work: 1.6456
@@
- id: 89, loan amount: 2100.0, should be paid back within 4 months
- id: 109, loan amount: 3139.45905, should be paid back within 5 months
[... and other loans.]
Demand deposit interest rate is 37.9503%, and loan rate is 40.7381%.
Your loanable amount is 265361.0053339647.
@@
- Education: The average price in the city is 65.7481, changed (8.2185%(up)) compared to last year. And the buildings price is 80.
@@
- Education: short of 905.9101880203923 units goods on the market
[... and other goods.]
During the past 12 months, the building has generated 18712.751 profit, running at an average efficiency of 0.424. Capital of the building has changed from 995.700 -> 1329.247 (333.547).
Last year, the building produced 24786.920 goods, and sold 502.607(2.0277%) goods. The storage has changed from {’Education’: 3679.6098, Health_Social_Work’: 4.052} to {’Education’: 24230.216, Public_Administration_Defence_SocialSecurity’: 5.984, Health_Social_Work’: 1.6456}
@@
- close_vacancy(position_id): Close a vacancy for a specific position. This action will set the salary of the position to 0, effectively closing the position temporarily.
@@
reasoning’: The building is currently facing efficiency issues, as it has no employees to utilize its capacity effectively. To improve productivity, we should increase salaries to attract skilled employees for essential positions. Simultaneously, as there is a shortage of educational goods in the market and the current price is higher than the average, we should consider adjusting prices to boost sales. Furthermore, considering the high demand and potential profit from education goods, we should also invest in durable goods to enhance production capacity.’,
@@
adjust_salary(2000067, 150)’,
adjust_salary(2000068, 150)’,
adjust_salary(2000069, 300)’,
capital_invest(500)’,
"adjust_price(’Education’, 85)"

### v2

@@ E.3 Full Examples
reasoning’: "Based on the city report, there are critical shortages in Electricity_Gas_Steam_AirConditioning, Telecommunications, IT Information Services, Construction, Motor Vehicles, Mining, Machinery, Real Estate, Manufacturing, Paper Printing, Food, Agricultural, Textiles, Financial, Chemical, Rubber/Plastics, Publishing, Fabricated Metals, Education, Health, Public Services, Entertainment, Aquaculture, Pharmaceuticals, Electrical Equipment, Transport, Logistics, and Service Activities. Many of these shortages are significant, particularly Construction, Motor Vehicles, Mining Products, Manufacturing, Real Estate, Energy, and Telecommunications. The residential sector has a high homeless rate even though there is reported vacant housing, indicating rent is likely unaffordable. Factory buildings should be clustered together and away from residential blocks per guidelines; most factories are concentrated on the left and center-bottom of the map. To address the most urgent needs and the highest lack,’ I will prioritize: 1) Energy Production Facility (for electricity), 2) TechHub (for IT services/telecommunications), 3) Vehicle Manufacturing Plant (motor vehicles), 4) Mining Facility, 5) Food Processing Plant, 6) Construction Firm, 7) Community Health Center, 8) Educational Institution, 9) Real Estate Development Office. Factory sites will be placed in zones with existing factories (to the west, southwest, and center), while public/community services (health, education, real estate) should be placed near residential/high-density business clusters (southeast, east, and center-right). Public services will be distributed near residential clusters to optimize access. Each location is selected for minimal disruption and adjacency to similar use, supporting the existing city layout logic.",
@@
{’x’: -5, y’: 12, type’: Energy Production Facility’},
{’x’: -6, y’: 13, type’: TechHub’},
{’x’: -6, y’: 10, type’: Vehicle Manufacturing Plant’},
{’x’: -5, y’: 8, type’: Mining Facility’},
{’x’: -4, y’: 12, type’: Food Processing Plant’}
@@
Here is the monthly report for Educational Institution (No. 200334):
Overall, during this month, the building gross profit is 6297.65, with rest cash 24053.69. It has 71.1184 durable investments for promoting production.
The building has 2 employees, 3 vacant positions, and run with efficiency 42.1176%.
@@
This month, the building income is 8676.186, which consists of the following parts:
- Local trading income: 8676.186(100.0000%)
This month, the building outcome is 2378.535, which consists of the following parts:
- Other outcome list as follows:
- salary: -2000.000(-84.0854%)
The efficiency of this building is 42.1176%, which consists of the following parts:
The building has 5 positions, and here are the details:
- position_id: 2003701, None(Maintenance Worker), salary: 350. has 29.4118% importance. But no one is working here. If an employee with average skill(Maintenance, Physical Labor) works here, about 388.5952 goods will be produced per month.
- position_id: 2003895, None(Administrative Assistant), salary: 250. has 23.5294% importance. But no one is working here. If an employee with average skill(Management) works here, about 315.4600 goods will be produced per month.
- position_id: 2003973, None(Janitor), salary: 300. has 11.7647% importance. But no one is working here. If an employee with average skill(Cleaning, Physical Labor) works here, about 162.9422 goods will be produced per month.
- position_id: 2000124, Ryan(Administrative Assistant), salary: 1700. has 23.5294% importance. Because of Ryans skills: Management*1.0 , the efficiency addition is 23.5294%.
- position_id: 2000440, William(Janitor), salary: 300. has 11.7647% importance. Because of Williams skills: Cleaning*0.98 Physical Labor*0.6 , the efficiency addition is 18.5882%.
If an employee with required skill lower than average (<1.0), you may decrease its salary moderately or fire him for a better ones. The unemployment rate in the city is 27.5833%, and the average unemployment rate in the past 12 months is 27.5833%.
You may adjust the salary of the position according to the unemployment rate.
If the unemployment rate is high, you could open more job positions (in practice, converting positions with zero wages into paid positions) while reducing the salaries of existing roles to save money.
Alternatively, if the unemployment rate is low, you may adjust the salary of the position to a higher level to attract more candidates. Additionally, you may want to close certain vacancies that consistently fail to attract applicants.
The building run with profit: 6297.6518. According to current dividend rate, 40.0000% of current balance will be distributed to shareholders, which is 9621.4769, and the rest 14432.2154 will stay in the building.
The social average return on investment is 1.3416%, and the interest rate of the bank is 0.6108%.
@@
- Accommodation_Food_Services: 0.0078
- Professional_Scientific_Technical_Activities: 0.0416
- Education: 0.2903
- Real_Estate: 0.0113
@@
- id: 8938, loan amount: 51443.301640482816, has been overdue for 2 months
Demand deposit interest rate is 0.6108%, and loan rate is 2.6108%.
Your loanable amount is 1984571.2363522195.
Here are the news in the past months:
1985-01-01 00:$00:$ A sudden collapse in food sector starts. All food related sectors production efficiency decreases to 50% of its original value.
@@
- Education: The average price in the city is 869.0021, changed (-10.6580%(down)) compared to last year. And the buildings price is 1035.4589.
@@
- Education: short of 29.099754566300394 units goods on the market
During the past 12 months, the building has generated 48954.522 profit, running at an average efficiency of 0.334. Capital of the building has changed from 74.265 -> 71.118 (-3.146).
Last year, the building produced 77.316 goods, and sold 76.925(99.4944%) goods. The storage has changed from {’Accommodation_Food_Services’: 0.0097, Education’: 0.4322742900636204, Professional_Scientific_Technical_Activities’: 0.020946254037423388, Public_Administration_Defence_SocialSecurity’: 0.017455211697852826, Wholesale_Retail_Trade_Vehicle_Repair’: 0.010909507311158015, Financial_Insurance_Activities’: 0.006545704386694809, Real_Estate’: 0.005672943801802168, Other_Service_Activities’: 0.005236563509355847, Administrative_Support_Services’: 0.005236563509355847, Health_Social_Work’: 0.004800183216909527} to {’Accommodation_Food_Services’: 0.0078, Professional_Scientific_Technical_Activities’: 0.0416, Education’: 0.2903, Real_Estate’: 0.0113}
@@
- remove_position(position_id): Remove a position from the building. If the position is occupied, the employee will be fired.
- add_position(): Add a new position to the building by sampling from the position templates.
@@
reasoning’: The building is currently experiencing a good profit margin and a surplus in cash, but it has a low efficiency rate of 64% primarily due to vacant positions. To optimize operations and maximize profit, new positions should be added, particularly for Civil Engineer and Site Foreman roles, which can significantly improve production efficiency. Furthermore, since the current unemployment rate is high, we can lower the salaries of existing roles to manage expenses while offering competitive salaries for new positions. We should also consider raising the price of Construction goods since there is a market shortage, which allows us to capitalize on this demand. Additionally, the building should focus on addressing overdue loans, potentially by borrowing more to clear outstanding debts if necessary.’,
@@
add_position()’,
add_position()’,
adjust_salary(2002754, 1500)’,
adjust_salary(2000007, 3500)’,
adjust_salary(2000083, 2500)’,
"adjust_price(’Construction’, 2000)",
borrow(30000)’